### v3

@@ 8 Devices
| Model | PixArt-XL-2-256-MS | FLUX.1[dev] | ||||||||
| Device Number | 1 | 4 | 8 | 2 | 4 | 8 | 1 | 4 | 8 | |
| w/ Orig. | - | 8.41 | 19.81 | 1.67 | 3.11 | 6.09 | - | 4.01 | 5.93 | |
FID Scores for Parallel Methods on the Pixart and Flux.1. w/ G.T. means calculating the FID metrics with the ground-truth images. w/ Orig. means calculating the FID metrics with the images generated by the 1-Device original implementation.
The maximum memory usage of sequence parallel and PipeFusion for SD3-medium and Flux.1-dev is depicted in Figure 10. The SD3-medium consists of a 2B parameter DiTs backbone (7.8 GB on disk) and text encoders of nearly 19GB on disk. The Flux.1-dev consists of a 12B parameter DiTs backbone (23 GB on disk) and text encoders of nearly 9.1GB on disk. The memory cost of PipeFusion is also significantly less than that of sequence parallelism. For the 12B Flux.1-dev, PipeFusion significantly reduces the memory footprint of model parameters. However, PipeFusion increases the memory usage of activations due to the consumption of the KV Buffer. Nevertheless, this increase is relatively minor compared to the overall model parameter memory usage. The overall memory usage of PipeFusion is 32% and 36% of SP on 1024px and 2048px cases using Flux.1.
5 Quality Results
In Figure 11, we show visual results. The images generated by PipeFusion are nearly indistinguishable from the original images to the human eye, across various configurations of the number of patches and devices. The table at the bottom of the figure includes the Fréchet Inception Distance (FID) Heusel et al. (2017), where a lower FID score is indicative of better performance, evaluated on Pixart and Flux.1. This data clearly indicates that PipeFusion outperforms DistriFusion in terms of FID across the same device configurations. This is because, compared to DistriFusion, PipeFusion utilizes more fresh KV values in its computations.
6 Ablation Study
@@
Warmup Step: The number of warmup steps can negatively impact the overall performance of PipeFusion. The input temporal redundancy is relatively small during the initial few diffusion timesteps. Therefore, it is usually to employ a few warmup steps to enable synchronous communication without resorting to stale activations, which would introduce waiting time, commonly referred to as bubbles, into the pipeline. As the number of sampling steps increases, the latency will also increase, as shown in Table 2. If the proportion of warmup steps is relatively small, no action is required. Our experiments involved a relatively small number of sampling steps, specifically 20 and 28. In other related work, it is common to use 50 steps Li et al. (2024) or even 100 steps Ma et al. (2024b), which helps to alleviate the overhead of warmup.
@@
7 Conclusion
This paper introduces PipeFusion to parallelize Diffusion Transformers (DiTs) inference on multiple devices. By exploiting the input similarities across diffusion timesteps and employing a patch-level pipelined approach to orchestrate communication and computation, PipeFusion reduces both communication bandwidth and memory demands, making DiT inference more efficient on PCIe-connected devices. I n the future, we plan to scale DiT inference to low-end multi-node GPU clusters, for example, PCIe inside a GPU node and Ethernet between nodes. We anticipate that PipeFusion will achieve even greater advantages in such environments.

### v4

@@ 8 Devices
Prompt: A pair of parking meters reflecting expired times.
Appendix B Ablation Study
@@
Warmup Step: The number of warmup steps can negatively impact the overall performance of PipeFusion. The input temporal redundancy is relatively small during the initial few diffusion timesteps. Therefore, it is usually to employ a few warmup steps to enable synchronous communication without resorting to stale activations, which would introduce waiting time, commonly referred to as bubbles, into the pipeline. As the number of sampling steps increases, the latency will also increase, as shown in Table 3. If the proportion of warmup steps is relatively small, no action is required. Our experiments involved a relatively small number of sampling steps, specifically 20 and 28. In other related work, it is common to use 50 steps [li2023distrifusion] or even 100 steps [ma2024deepcache], which helps to alleviate the overhead of warmup.
@@
| Method | 1024px | 2048px | 4096px |
| Tensor Parallel | 1.22 | 7.07 | 36.33 |
| DistriFusion | 0.77 | 2.48 | 25.24 |
| SP-Ring | 1.37 | 2.46 | 23.31 |
| SP-Ulysses | 2.59 | 3.54 | 27.41 |
| PipeFusion | 0.66 | 2.59 | 22.39 |
Evaluation on NVLink connected GPUs: To evaluate the sensitivity of PipeFusion to hardware interconnect bandwidth, we conducted additional experiments on 8A100 GPUs connected via NVLink (high-bandwidth) in contrast to the default 8L40 PCIe setting. Table 4 reports the latency for PixArt inference across different resolutions. PipeFusion consistently achieves the lowest latency across all resolutions and remains robust even under high-bandwidth NVLink settings. While the relative performance gap narrows at higher resolutions due to the increasing dominance of computation cost (i.e., the shift from communication- to computation-bound regimes), PipeFusion is still competitive or superior to all baselines. We also emphasize that the FID score is invariant to GPU type or interconnect bandwidth, as it is solely determined by the model architecture and inference algorithm.